# An Analytic Characterization of the Limb Asymmetry - Transit Time Degeneracy

### v1

@@ 5.3 Additional effects on the observability of limb asymmetry
Some exoplanets exhibit transit timing variations (TTVs) induced by gravitational interaction with other planets in the same system (e.g. Maciejewski et al., 2010; Adams et al., 2011). The amplitude of these TTVs can range from less than one second to tens of minutes depending on the orbital architecture of the system, and have variation periods upwards of hundreds of years. For certain planets, these TTVs may place a limit on the achievable precision on a planet’s time of conjunction that is comparable or worse than the timing biases derived here. Therefore, unless the underlying TTV trend is known extremely well, it will be difficult to disentangle signatures of limb asymmetry from the uncertainty in the time of conjunction generated by TTV interactions for such planets.
@@
Finally, our analysis constructed the disk of an asymmetric-limb planet by explicitly assuming east-west asymmetry where each hemisphere is a uniform semi-circle. Real exoplanet atmospheres likely do not follow this geometry. For example, strong meridional circulation that preferentially transports condensates to a planet’s polar regions may cause pole-equator asymmetry that is more significant than east-west asymmetry (e.g. see Parmentier et al., 2013; Charnay et al., 2015; Line & Parmentier, 2016). In this case, the planetary disk would be shaped differently than we have assumed, either as a 90-degree rotation or even a non-semi-circular shape, which may lead to a different form of the limb asymmetry - transit timing degeneracy than we have derived here.
@@ 5.4 Potential Limitations of Our Analytic Model
Our analytic model neglects any rotation of the planetary disk during the transit. For a planet with asymmetric-limbs, especially as constructed herein, the projected planetary disk will change slightly as the planet rotates slightly between ingress and egress. We neglected this rotation to keep the geometry as simple as possible, which will cause our formulae to slightly underestimate the timing bias, for example contributing to the outstanding residual in Figure 4. Fortunately, for the planets we considered here, which have semi-major axes ranging $\sim$9 - 18x their host stellar radii, this is a negligible effect. To test the maximum effect this will have, we repeated our calculations for the exoplanet WASP-12b, which has $a/R_{\star}=3$ (Chakrabarty & Sengupta, 2019) and represents the population for which intratransit rotation is most important. In this case, our analytic model underestimated the timing bias by 0.39 s for N=1, 0.78 s for N=2, and 1.93 s for the extreme case of N=5.
@@
Similarly, in our analytic model, we assume the planets move in a flat plane along the stellar disk during transit. In reality, of course, orbits are curved so these paths are slightly non-linear, especially for close-in orbits as above. To examine the validity of this assumption, we re-derived the timing bias based on the transit duration derived by Seager & Mallén-Ornelas (2003) which solves exactly for the sky-projected planet-star separation with time, and compared it to our calculations for WASP-39b and WASP-12b. The latter case, as before, represents the most close-in orbits possible where this approximation would break down. In both cases though, the difference relative to our nominal value is effectively zero.
@@
As mentioned previously, the semi-circle based model that we use for our asymmetric-limb planetary disk, both when deriving our analytic model and when using catwoman in our numerical model, is not strictly realistic. As seen clearly in the exaggerated example in Figure 1, this construction creates a discontinuity in the atmosphere near the planet’s poles, which becomes more extreme for greater limb asymmetry. This limitation, however, is unlikely to be important given current and near-future data quality. For example, as seen in the residuals in Figure 4 and discussed in Section 4.1, this discontinuity only induces error in our model of approximately one tenth of a second for limb asymmetry greater than five scale heights. While this was calculated just for one example exoplanet, the magnitude would not be significantly different for other planets. Transit timing at such high precision is likely not achievable with current instrumentation, as combined fits of state-of-the-art JWST, TESS, Spitzer, and ground-based data have only been able to achieve 1-$\sigma$ uncertainty on $t_{c}$ of $\sim$0.6 s or worse (Carter et al., 2024; Murphy et al., 2024). Also, it is not clear from current observations whether limb asymmetry as large as five scale heights and greater actually exists in nature. Nevertheless, analytic and numerical models which define fully continuous atmospheric surfaces, such as the transmission string numerical model (Grant & Wakeford, 2023), may improve upon our result.
@@ 6 Summary
Measuring morning-to-evening variations of temperature, elemental and molecular abundances, and aerosol properties in exoplanet atmospheres is critical to understanding the underlying dynamics and chemistry of the atmosphere. Transit observations are able to extract the separate contributions from a planet’s morning and evening terminator using high-resolution ground-based spectroscopy and, more recently, low-resolution space-based spectroscopy with JWST. A strong degeneracy exists between the effect of limb asymmetry on a light curve and uncertainty in the planet’s time of conjunction. In this work, we discussed the origin of this degeneracy and derived a simple, physically-motivated analytical model for it, highlighted by Equation 2 and generalized in Equation 14. We found that limb asymmetry of one to two scale heights difference in limb radii can cause the apparent time of conjunction to be biased by upwards of ten seconds. This bias is more extreme for planets with larger scale height, and as mentioned we derived a generalized formula (Equation 14) for estimating this bias on any exoplanet. When comparing our analytic model to simulated observations of an asymmetric-limb planet fit using a uniform-limb model, we found an additional numerical bias is added to to the inferred time of conjunction, which is typically less than one second. We dubbed this the “extra-area compensation effect” as it stems from the relative occulting areas of a uniform-limb disk and asymmetric-limb disk when stacked atop one another. This effect is included in the above generalized formula, and we derived formulae (Equation 12, generalized in Equation 15) specifically for this effect as well.
Our results underscore the importance of obtaining precise $t_{c}$ measurements of target exoplanets to be able to probe for limb asymmetry in their atmospheres. We have shown that times derived from single-filter observations can be strongly biased. Therefore, the recommended method to make precise $t_{c}$ measurements is to simultaneously fit transit observations, at as high cadence and precision as possible, from multiple wavelengths, as shown by Powell et al. (2019). Optical space-based observations (e.g., TESS, CHEOPS, Kepler) are usually readily available, but have a lot of bandpass overlap. We would recommend combining these with non-overlapping ground-based and/or near-infrared space-based observations (e.g., Spitzer/IRAC, and JWST) to get the best possible constraint. These observations need not be simultaneous with each other or the primary observation (i.e., the one being fitted for limb asymmetry), but will be most effective when as close in time to the primary observation as possible.

### v2

@@ 5.3 Additional Effects on the Observability of Limb Asymmetry
Some exoplanets exhibit transit timing variations (TTVs) induced by gravitational interaction with other planets in the same system (e.g. Maciejewski et al., 2010; Adams et al., 2011). The amplitude of these TTVs can range from less than one second to tens of minutes, depending on the orbital architecture of the system, and have variation periods upward of hundreds of years. For certain planets, these TTVs may place a limit on the achievable precision on a planet’s time of conjunction that is comparable or worse than the timing biases derived here. Therefore, unless the underlying TTV trend is known extremely well, it will be difficult to disentangle signatures of limb asymmetry from the uncertainty in the time of conjunction generated by TTV interactions for such planets.
@@
Finally, our analysis constructed the disk of an asymmetric-limb planet by explicitly assuming east–west asymmetry where each hemisphere is a uniform semicircle. Real exoplanet atmospheres likely do not follow this geometry. For example, strong meridional circulation that preferentially transports condensates to a planet’s polar regions may cause pole–equator asymmetry that is more significant than east–west asymmetry (e.g. see Parmentier et al., 2013; Charnay et al., 2015; Line & Parmentier, 2016). In this case, the planetary disk would be shaped differently than we have assumed, either as a 90 rotation or even a non-semicircular shape, which may lead to a different form of the limb asymmetry - transit timing degeneracy than we have derived here.
@@ 5.4 Potential Limitations of Our Analytic Model
Our analytic model neglects any rotation of the planetary disk during the transit. For a planet with asymmetric limbs, especially as constructed herein, the projected planetary disk will change slightly as the planet rotates slightly between ingress and egress. We neglected this rotation to keep the geometry as simple as possible, which will cause our formulae to slightly underestimate the timing bias – for example, contributing to the outstanding residual in Figure 4. Fortunately, for the planets we considered here, which have semi-major axes ranging $\sim$9 – 18x their host stellar radii, this is a negligible effect. To test the maximum effect this will have, we repeated our calculations for the exoplanet WASP-12b, which has $a/R_{\star}=3$ (Chakrabarty & Sengupta, 2019) and represents the population for which intratransit rotation is most important. In this case, our analytic model underestimated the timing bias by 0.39 s for $N$=1, 0.78 s for $N$=2, and 1.93 s for the extreme case of $N$=5.
@@
Similarly, in our analytic model, we assume the planets move in a flat plane along the stellar disk during transit. In reality, of course, orbits are curved so these paths are slightly nonlinear, especially for close-in orbits as above. To examine the validity of this assumption, we rederived the timing bias based on the transit duration derived by Seager & Mallén-Ornelas (2003) which solves exactly for the sky-projected planet–star separation with time, and compared it to our calculations for WASP-39b and WASP-12b. The latter case, as before, represents the most close-in orbits possible where this approximation would break down. In both cases, though, the difference relative to our nominal value is effectively zero.
@@
As mentioned previously, the semicircle-based model that we use for our asymmetric-limb planetary disk, both when deriving our analytic model and when using catwoman in our numerical model, is not strictly realistic. As seen clearly in the exaggerated example in Figure 1, this construction creates a discontinuity in the atmosphere near the planet’s poles, which becomes more extreme for greater limb asymmetry. This limitation, however, is unlikely to be important given current and near-future data quality. For example, as seen in the residuals in Figure 4 and discussed in Section 4.1, this discontinuity only induces error in our model of approximately one tenth of a second for limb asymmetry greater than five scale heights. While this was calculated just for one example exoplanet, the magnitude would not be significantly different for other planets. Transit timing at such high precision is likely not achievable with current instrumentation, as combined fits of state-of-the-art JWST, TESS, Spitzer, and ground-based data have only been able to achieve 1-$\sigma$ uncertainty on $t_{c}$ of $\sim$0.6 s or worse (Carter et al., 2024; Murphy et al., 2024). Also, it is not clear from current observations whether limb asymmetry as large as five scale heights and greater actually exists in nature. Nevertheless, analytic and numerical models that define fully continuous atmospheric surfaces, such as the transmission string numerical model (Grant & Wakeford, 2023), may improve upon our result.
@@ 6 Summary
Measuring morning-to-evening variations of temperature, elemental and molecular abundances, and aerosol properties in exoplanet atmospheres is critical to understanding the underlying dynamics and chemistry of the atmosphere. Transit observations are able to extract the separate contributions from a planet’s morning and evening terminator using high-resolution ground-based spectroscopy and, more recently, low-resolution space-based spectroscopy with JWST. A strong degeneracy exists between the effect of limb asymmetry on a light curve and uncertainty in the planet’s time of conjunction. In this work, we discussed the origin of this degeneracy and derived a simple, physically motivated analytical model for it, highlighted by Equation 2 and generalized in Equation 14. We found that limb asymmetry of one to two scale heights difference in limb radii can cause the apparent time of conjunction to be biased by upward of 10 seconds. This bias is more extreme for planets with larger scale height, and as mentioned we derived a generalized formula (Equation 14) for estimating this bias on any exoplanet. When comparing our analytic model to simulated observations of an asymmetric-limb planet fit using a uniform-limb model, we found an additional numerical bias is added to to the inferred time of conjunction, which is typically less than one second. We dubbed this the “extra-area compensation effect,” as it stems from the relative occulting areas of a uniform-limb disk and asymmetric-limb disk when stacked atop one another. This effect is included in the above generalized formula, and we derived formulae (Equation 12, generalized in Equation 15) specifically for this effect as well.
Our results underscore the importance of obtaining precise $t_{c}$ measurements of target exoplanets to be able to probe for limb asymmetry in their atmospheres. We have shown that times derived from single-filter observations can be strongly biased. Therefore, the recommended method to make precise $t_{c}$ measurements is to simultaneously fit transit observations, at the highest possible cadence and precision, from multiple wavelengths, as shown by Powell et al. (2019). Optical space-based observations (e.g., TESS, CHEOPS, and Kepler) are usually readily available, but have a lot of bandpass overlap. We would recommend combining these with nonoverlapping ground-based and/or near-infrared space-based observations (e.g., Spitzer/IRAC, and JWST) to get the best possible constraint. These observations need not be simultaneous with each other or the primary observation (i.e., the one being fitted for limb asymmetry), but will be most effective when as close in time to the primary observation as possible.